OpenStudy (anonymous):

1. While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3?

OpenStudy (anonymous):

Is this supposed to be: Peter wrote on his paper that (x^3)*(x^3)*(x^3)*(x^3) equaled x^(3+3+3+3)? If so, then Peter did simplify his work correctly and completely. And no, Peter's work would not be the same if he were to simply (x^3) + (x^3) + (x^3) + (x^3): (x^3)*(x^3)*(x^3)*(x^3) = x^(3+3+3+3) = x^12 whereas, (x^3) + (x^3) + (x^3) + (x^3) = 4*(x^3)

OpenStudy (anonymous):

The information included in this question is the intellectual property of Florida Virtual School. Posting questions from any FLVS course and receiving answers is considered cheating. When you are having help with your work, all you need to do is call your instructor - they are there to help you every step of the way!